### v1

# Learning with Noisy Labels: the Exploration of Error Bounds in Classification

Abstract
Numerous studies have shown that label noise can lead to poor generalization performance, negatively affecting classification accuracy. Therefore, understanding the effectiveness of classifiers trained using deep neural networks in the presence of noisy labels is of considerable practical significance. In this paper, we focus on the error bounds of excess risks for classification problems with noisy labels within deep learning frameworks. We begin by exploring loss functions with noise-tolerant properties, ensuring that the empirical minimizer on noisy data aligns with that on the true data. Next, we estimate the error bounds of the excess risks, expressed as a sum of statistical error and approximation error. We estimate the statistical error on a dependent (mixing) sequence, bounding it with the help of the associated independent block sequence. For the approximation error, we first express the classifiers as the composition of the softmax function and a continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The main task is then to estimate the approximation error for the continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. Finally, we focus on the curse of dimensionality based on the low-dimensional manifold assumption.
1 Introduction
Deep learning has made breakthroughs in the fields of computer vision [SLD17], natural language processing [YHPC18], medical image processing [SK22] and robotics [LTZ+22], which proves its strong efficiency in solving complicated problems. Leveraging high-performance computing resources and high-quality data, neural network-based methods outperform traditional machine learning approaches across various applications, including image classification, object detection, and speech recognition.
Over the years, the application of deep neural networks to classification problems has garnered significant attention. In practical applications, the data we collect may be disturbed by noise. Literature typically classifies noise into two types: attribute noise and class noise [ZW04]. Attribute noise refers to disturbances that affect feature values, such as the addition of small Gaussian noise during measurement. In contrast, class noise impacts the labels assigned to instances, directly affecting the sample labels rather than the features. Some studies have indicated that random class noise can be more detrimental than random attribute noise, highlighting the importance of addressing class noise [SGLH14].
Labeling large-scale datasets is a costly and error-prone process, and even high-quality datasets are inevitably subject to mislabeling errors. Labels that may contain errors are referred to as noisy labels, which can arise from various sources. Due to the high cost of labeling, some labels are obtained from non-expert sources on websites, such as image search engines or web crawlers. Even data labeled by experts may contain errors; these classification mistakes are not solely due to human oversight but can also result from the widespread use of automated classification systems. Additionally, the limitations of descriptive language can contribute to the occurrence of noisy labels [BC90]. Furthermore, some artificial noises are introduced to protect individuals’ privacy [vdHvdH02, GGK+21].
Numerous studies have demonstrated that label noise can lead to poor generalization performance, negatively impacting classification accuracy [FV14, NOPF10]. Therefore, the effectiveness of classifiers trained using deep neural networks in the presence of noisy labels is of considerable practical importance.
Although neural networks have achieved remarkable results in real applications, our theoretical understanding of them remains limited. The expressive power of neural networks refers to their ability to approximate various functions. Understanding the approximation capabilities of deep neural networks has become a fundamental research direction for uncovering the advantages of deep learning over traditional methods, with related studies dating back to at least the 1980s. The Universal Approximation Theorem [Bar94] states that a neural network with an appropriate activation function and a depth of two can approximate any continuous function on a compact domain to any desired precision. This illustrates the powerful expressive capability of neural networks and provides a theoretical foundation for their application. However, in practice, this approximation theorem has significant limitations. For instance, approximating complex functions may require a large number of neurons, leading to computational difficulties.
Currently, most analyses of excess risk error bounds are focused on regression models, while research on error bounds for classification problems within deep learning frameworks remains relatively scarce, especially in the context of finite samples and label noise. This paper aims to fill this gap. We begin with noise-tolerant property of the loss functions, which demonstrates that the empirical minimizer on noisy data aligns with that on the true data. To estimate the error bound, we specifically divide them into approximation error and statistical error, which can be estimated separately. Our contributions can be summarized as follows:
(1) To the best of our knowledge, we are the first to provide an error bound for the classification problem with noisy label. The main results are in Theorems 4.1 and 6.1.
(2) The noise-tolerant property of the loss function: The noise-tolerant property guarantees the relationship of the empirical minimizers on both noisy data and true data. It is important to note that both the $\ell_{1}$ loss and the reverse cross entropy (RCE) loss exhibit noise-tolerant properties. However, for the cross entropy (CE) loss, this noise-tolerant property holds only when the classifier $f$ is a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$, where $K$ represents the number of classes and $e_{i}$ is the unitary vector whose $i$-th component is one, with all other components being zero.
(3) The statistical error: In practice, the samples in the dataset may not be independent. In this paper, we focus on a dependent (mixing) sequence. We bound the statistical error with the help of the associated independent block (IB) sequence.
(4) The approximation error: The classifier we aim to develop is designed to map input features to a $K$-dimensional unitary vector. To achieve this, we first express the classifiers as a composition of the softmax function and a continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The primary objective is to estimate the approximation error for this continuous function. Although our theory emphasizes $K$-dimensional unitary outputs, it remains valid when we change the output space to $\mathcal{E}$, provided that the outputs of the true classifier and the optimal classifier are sufficiently separated from the margins.
(5) The curse of dimensionality: Based on the low-dimensional manifold assumption, we effectively alleviate the issue of the curse of dimensionality.
The rest of this paper is organized as follows. Section 2 reviews the related works about noisy labels and error estimations. Section 3 presents the problem setup, including the definitions of the expected excess risk estimations on the true and noisy distributions, the empirical excess risk estimations on the true and noisy data, the neural networks we considered in this paper, which is followed by the theoretical results of the excess risk under noisy labels in Section 4. Some technical details of proof are provided in Section 5. In Section 6, we present a theoretical result aimed at alleviating the curse of dimensionality. Lastly, we conclude the results in Section 7.
2 Related Works
2.1 Noisy label
Learning from noisy-labeled data is an active area of research. Learning discriminative models was formulated using undirected graphical models [Vah17] or probabilistic graphical models [XXY+15], which represent the relationship between noisy and clean labels in a semi-supervised setting. Sukhbaatar et al. [SBP+15] introduced an additional noise layer in the network that adapted the outputs to align with the noisy label distribution. Patrini et al. [PRKM+17] proposed two procedures for loss correction that were agnostic to both the application domain and the network architecture.
Furthermore, decoupling methods were designed for training with noisy labels using two models. Malach et al. [MSS17] introduced a meta-algorithm to tackle the issue of noisy labels by training two networks and establishing update rules that decoupled the decisions of ‘when to update’ and ‘how to update’. Following this, the Co-teaching algorithm [HYY+18] was developed, enabling two deep neural networks to train concurrently while teaching each other under mini-batch conditions. Building on the Decoupling and Co-teaching algorithms, the Co-teaching+ algorithm [YHY+19] was introduced. By implementing the strategy of ‘Update by Disagreement’, this method prevented the two networks from becoming overly consistent during training.
Additionally, Lee et al. [LHZY18] addressed noisy labels through transfer learning methods, introducing CleanNet, a joint neural embedding network that provided insights into label noise by manually verifying only a small subset of classes and passing this knowledge to other classes. Jiang et al. [JZL+18] proposed an algorithm for joint optimization of large-scale data using MentorNet, which supervised the training of a basic deep network, StudentNet, enhancing its generalization ability when trained on noisy labels.
Xia et al. [XLH+20] introduced a robust early learning method that distinguished between critical and non-critical parameters, applying different update rules to gradually reduce non-critical parameters to zero, thereby suppressing the impact of noisy labels. Li et al. [LXGL22] proposed selective supervised contrastive learning (Sel-CL), which trained with confident and noisy pairs, enabling the model to distinguish between noise and mitigate its influence on supervised contrastive learning (Sup-CL). An alternative approach was to seek loss functions that were inherently noise-tolerant [GKS17]. The authors provided sufficient conditions for a loss function to ensure that risk minimization under that function was inherently resilient to label noise in multiclass classification problems. Subsequently, symmetric cross-entropy was introduced for robust learning with noisy labels [WMC+19]. Wei et al. [WFS+23] presented a framework for label noise in fine-grained datasets, stochastic noise-tolerated supervised contrastive learning (SNSCL), which mitigated the effects of noisy labels by designing noise-tolerant loss functions and incorporating random modules.
2.2 Error estimation
In recent years, research on the expressive and approximation capabilities of neural networks has become increasingly active. The approximation error of feedforward neural networks with various activation functions has been studied for different types of functions. Lu et al. [LPW+17] examined the influence of width on the expressive ability of neural networks, establishing corresponding approximation error estimates for Lebesgue integrable functions. They obtained a general approximation theorem for ReLU networks with bounded width, demonstrating that a ReLU network with a width of $d+4$ is a universal approximator, where $d$ is the input dimension. Lu et al. [LSYZ21] established the approximation error bounds of ReLU neural networks for smoothing functions concerning width and depth. They proved that a deep neural network with width $O(N\ln N)$ and depth $O(L\ln L)$ can approximate functions $f\in C^{s}([0,1]^{d})$, with an approximation error of $O\left(\left\|f\right\|_{C^{s}([0,1]^{d})}N^{-2s/d}L^{-2s/d}\right)$ for any $N,L\in\mathbb{N}^{+}$, where $\left\|f\right\|_{C^{s}([0,1]^{d})}:=\max\left\{\left\|\partial^{\alpha}f\right\|_{L^{\infty}([0,1]^{d})}:\left\|\alpha\right\|_{1}\leq s,\alpha\in\mathbb{N}^{d}\right\}$. Petersen et al. [PV18] established the best approximation error bounds for approximating piecewise $C^{\beta}$ functions on the interval $[-1/2,1/2]^{d}$ using ReLU neural networks. Hadrien et al. [MYD21] provided an error estimate for approximating generalized finite-bandwidth multivariate functions with deep ReLU networks. Collectively, these studies focus on the ability of ReLU-activated feedforward neural networks to approximate various types of functions.
Additionally, many studies have explored other activation functions. Danilo et al. [DC13] investigated the approximation order of neural network operators with sigmoid activation functions using a moment-type approach. DeRyck et al. [DLM21] provided error bounds for neural networks utilizing hyperbolic tangent activation functions in the approximation of Sobolev regular and analytic functions, proving that tanh networks with only two hidden layers can approximate functions at a faster or equivalent rate compared to deeper ReLU networks.
Another direction of research focuses on the error bounds of regression model in $\beta$-Hölder smooth space, which are established to evaluate approximation quality.
Stone [Sto82] assumed that the true regression model is $\beta$-Hölder smooth, where the smoothness index $\beta>0$. The convergence rate of the prediction error is $C_{d}n^{-2\beta/(2\beta+d)}$. This indicates that the convergence rate is exponentially dependent on the dimension $d$ of the inputs, while the prefactor $C_{d}$ is independent of the sample size $n$ but depends on $d$ and other model parameters. Research by Schmidt-Hieber [SH20] and Nakada and Imaizumi [NI20] demonstrated that deep neural network regression can achieve the optimal minimax rate established by Stone in 1982 under certain conditions. Shen, Yang and Zhang [SYZ20] derived the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\leq 1$, finding that the prefactor depends on a polynomial of $d$. Lu et al. [LSYZ21] studied the approximate error bound for $\beta$-Hölder continuous functions when the smoothness index $\beta\geq 1$, noting that its prefactor relates to an exponential form of $d$.
Jiao et al. [JSLH23] improved the result of [LSYZ21]. For the case of the smoothness index $\beta>1$, they found that the non-asymptotic upper limit of the prediction error of the empirical risk minimizer is polynomially dependent on the dimension $d$ of the predictor rather than exponentially dependent. Chen et al. [CJLZ22] focused on nonparametric regression of Hölder functions on low-dimensional manifolds using deep ReLU networks. They proved the adaptability of deep ReLU networks to low-dimensional geometric structures in data, partially explaining their effectiveness in handling high-dimensional data with low-dimensional geometric structures. Feng et al. [FJK+23] recently addressed over-parameterized deep nonparametric regression. By effectively balancing approximation and generalization errors, they derived results for Hölder functions with constrained weights, highlighting their application in reinforcement learning.
Furthermore, numerous studies have concentrated on the curse of dimensionality, a phenomenon characterized by a significantly slowed convergence rate when the input dimension $d$ is excessively large. This challenge is inherent in high-dimensional computations, rendering it a pivotal area of investigation.
Yang et al. [MYD21] focused on constraining the objective function space and demonstrated a theorem on the approximation of generalized bandlimited multivariate functions using deep ReLU networks, which overcame the curse of dimensionality. This theorem builds upon Maurey’s work and leverages the proficiency of deep ReLU networks in efficiently approximating Chebyshev polynomials and analytic functions. Shen et al. [SYZ21] considered neural networks equipped with alternative activation functions introduced Floor-ReLU networks, achieving root exponential convergence while simultaneously avoiding the curse of dimensionality for (Hölder) continuous functions. These networks provide an explicit error bound in deep network approximation. Jiao et al. [JSLH23] was grounded in the low-dimensional subspace assumption and illustrated that the neural regression estimator could bypass the curse of dimensionality under the premise that the predictor resides on an approximate low-dimensional manifold or a set possessing low Minkowski dimension.
3 Setup of Classification Problem with Noisy Label
We consider a $K$-class classification problem with noisy label. We start from the expected risk, which quantifies the average loss of a classifier over the entire population or distribution of data.
Definition 3.1.
Let $Z=(X,Y,Y^{\eta})$ be a tuple of random vectors, where $(X,Y)$ follow a distribution $D$, $X\in{\mathbb{R}}^{d}$ represents the feature vector, $Y,Y^{\eta}\in\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}\in{\mathbb{R}}^{K}|y_{i}\in[0,1],\sum^{K}_{i=1}y_{i}=1\}$ denote the true label and the random noisy vector of $Y$, respectively. Let $f$ be a map which is from ${\mathbb{R}}^{d}$ to $\mathcal{Y}$ and $\ell(\cdot,\cdot)$ be a loss function. Then the expected risks associated with true distribution or noisy distribution are defined as
$\mathcal{L}(f)=\mathbb{E}_{X,Y}[\ell(f(X),Y)],\quad\mathcal{L}^{\eta}(f)=\mathbb{E}_{X,Y^{\eta}}[\ell(f(X),Y^{\eta})].$ (3.1)
In practice, the distribution $D$ and its perturbed versions are often unknown. Consequently, we typically focus on exploring the empirical risks at the sample level.
Definition 3.2.
Let $(x,y,y^{\eta})=\{(x_{i},y_{i},y^{\eta}_{i})\}_{i=1}^{n}$ be the dataset, where $x_{i},y_{i},y^{\eta}_{i}$ represent the feature, true label, and noisy label of the $i$-th data point, respectively. Let $f$ be a map from ${\mathbb{R}}^{d}$ to $\mathcal{Y}$, and let $\ell(\cdot,\cdot)$ denote a loss function. Then the empirical risks about the true data or the noisy data are
$\mathcal{L}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i}),y_{i}),\quad\mathcal{L}^{\eta}_{n}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell(f(x_{i}),y^{\eta}_{i}).$ (3.2)
A loss function, denoted $\ell(\cdot,\cdot)$, is commonly introduced to quantify the differences between $f(x)$ and its associated label $y$. This includes various loss functions such as the $\ell_{1}$ loss, the $\ell_{2}$ loss, the CE loss and the RCE loss [WMC+19], among others.
In practice, it is infeasible to explore all possible functions to find the optimal classifier. In recent years, a series of research results on deep neural networks have been achieved with great success. We select a class of deep neural networks with a specific structure. Let ${\mathcal{F}}_{d,K}(\mathcal{W},\mathcal{D})$ be a class of a deep neural network with an input dimension of $d$, an output dimension of $K$, width ${\mathcal{W}}$ and depth ${\mathcal{D}}$. In the following, we use $\mathcal{F}$ as a shorthand when the parameters used in the neural networks do not need to be emphasized.
In this paper, we specifically consider a class of the ReLU neural networks. Let $\phi(\cdot)\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ be a function that maps $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The function $\phi$ can be constructed as follows:
$\phi_{0}(x)=x,\ \phi_{i+1}(x)=\sigma(W_{i}\phi_{i}(x)+b_{i}),\ i=0,1,\cdots,\mathcal{D}-1,\ \phi(x)=W_{\mathcal{D}}\phi_{\mathcal{D}}(x)+b_{\mathcal{D}}.$
The weight matrix is given by $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$, and the bias vector is $b_{i}\in\mathbb{R}^{p_{i+1}}$, where $p_{i}$ is the width of the $i$-th layer for $i=0,\cdots,{\mathcal{D}}$, the activation function is defined as $\sigma(x)=\max(x,0)$. Let $\mathrm{softmax}:{\mathbb{R}}^{K}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}\in{\mathbb{R}}^{K}|y_{i}\in[0,1],\sum^{K}_{i=1}y_{i}=1\}$ be the softmax function, then $\mathrm{softmax}\circ\phi:[0,1]^{d}\to\mathcal{Y}$.
The structure of the neural network above consists of $\mathcal{D}$ hidden layers and a total of $(\mathcal{D}+2)$ layers, with $p_{0}=d$ and $p_{\mathcal{D}+1}=K$. The width of the neural network satisfies $\mathcal{W}=\max\{p_{1},\cdots,p_{{\mathcal{D}}}\}$, the size of the neural network is given by $\mathcal{Q}=\sum_{i=0}^{\mathcal{D}}p_{i+1}\times(p_{i}+1)$, and the number of neurons is $\mathcal{U}=\sum_{i=1}^{\mathcal{D}}p_{i}$.
As a subset of ${\mathcal{F}}_{d,K}(\mathcal{W},\mathcal{D})$, a class of the norm-constrained neural network, denoted as $\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D},\mathcal{M})$, represents the collection of functions $\phi(\cdot)\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ subject to the constraint:
$\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|\prod_{i=0}^{{\mathcal{D}}-1}\max\left\{\|(W_{i},b_{i})\|,1\right\}\leq{\mathcal{M}},$
[…]
Throughout the paper, we use $f_{0}:=\arg\min_{f}\mathcal{L}(f)$, $f^{\eta}_{0}:=\arg\min_{f}\mathcal{L}^{\eta}(f)$ as the global minimizers of the expected risks of the true distribution and the noisy distribution. Similarly, we define
$\hat{f_{n}}=\mathop{\arg\min}_{f\in{\mathcal{F}}}\mathcal{L}_{n}(f),\ \hat{f^{\eta}_{n}}=\mathop{\arg\min}_{f\in{\mathcal{F}}}\mathcal{L}^{\eta}_{n}(f),$ (3.3)
as the empirical risk minimizer for the true data and the noisy data, where ${\mathcal{F}}$ represents a deep neural network class with a certain structure. Additionally, $\lfloor x\rfloor$ denotes the largest integer no greater than $x$, $\lceil x\rceil$ denotes the smallest integer no less than $x$, and $[J]=\{0,\cdots,J-1\},\ \forall J\in\mathbb{N}^{+}$.
4 The Excess Risk under Noisy Label
We consider the following classification problem
$Y^{\eta}=n_{\eta}(Y)=n_{\eta}(f_{0}(X)),$
where $X$ represents the features, $Y,Y^{\eta}$ are the true and noisy labels, respectively, $f_{0}(\cdot)$ is the unknown classifier, $n_{\eta}(\cdot)$ is the disturbing function which maps the true label to the noisy label. We assume symmetric or uniform noise. Let $\eta\in[0,1]$ denote the noise rate. For all $i,j\in\{1,\cdots,K\}$, $Y^{\eta}$ obeys the following conditional probability conditioned on $Y$
$\mathrm{Prob}(Y^{\eta}=e_{i}|Y=e_{j})=\left\{\begin{aligned} &1-\eta,&&i=j,\\ &\frac{\eta}{K-1},&&i\neq j.\end{aligned}\right.$ (4.1)
In this section, we explore the theoretical bounds about classification problem with noisy labels. We focus on the excess risks on the true and noisy distributions, which are defined as
$\begin{split}\mathcal{L}(\hat{f}^{\eta}_{n})-\mathcal{L}(f_{0})=&\mathbb{E}_{X,Y}[\ell(\hat{f}^{\eta}_{n}(X),Y)]-\mathbb{E}_{X,Y}[\ell(f_{0}(X),Y)],\\ \mathcal{L}^{\eta}(\hat{f}^{\eta}_{n})-\mathcal{L}^{\eta}(f_{0})=&\mathbb{E}_{X,Y^{\eta}}[\ell(\hat{f}^{\eta}_{n}(X),Y^{\eta})]-\mathbb{E}_{X,Y^{\eta}}[\ell(f_{0}(X),Y^{\eta})].\end{split}$
We begin with an assumption regarding the logarithm function and then provide the bounds on the excess risks.
Assumption 4.1.
Let $\epsilon>0$ be a small positive number, we define
$\log x=\left\{\begin{array}[]{ll}\log x,&x\geq\epsilon,\\ A=\log\epsilon,&x\in[0,\epsilon).\end{array}\right.$
This technique in Assumption 4.1 is reasonable, which is similar to the clipping operation implemented by most deep learning frameworks.
Theorem 4.1.
Let $\kappa$ be a continuous map from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$ and define $f_{0}=\mathrm{softmax}\circ\kappa:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0\ \forall i,\sum^{K}_{i=1}y_{i}=1\}$. Let ${\mathcal{F}}_{d,K}(\mathcal{W},\mathcal{D},{\mathcal{M}})$ be a class of a norm-constrained deep neural network with an input dimension $d$, an output dimension $K$, width ${\mathcal{W}}$, depth ${\mathcal{D}}$, and ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\{\|(W_{i},b_{i})\|_{1,\infty},1\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer. We consider a strictly stationary $\beta$-mixing sequence defined in Definition 5.3. Let $n$ be the sample size and $a_{n},n/a_{n}$ be positive integers. For any $N,M\in\mathbb{N}^{+}$, when ${\mathcal{W}}=\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\}$, ${\mathcal{D}}=11M+18$ and Assumption 4.1 holds, the excess risks for the empirical risk minimizer $\hat{f^{\eta}_{n}}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ can be bounded as
(1) when the loss we consider is either the $\ell_{1}$ or RCE loss
$|\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})|\leq$ $\frac{8\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda Kn\beta_{a_{n}}}{a_{n}}$
$\qquad+131c\lambda\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$ (4.2)
(2) when the loss function we consider is one of the $\ell_{1}$, CE and RCE losses
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq$ $\frac{8\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{4\lambda Kn\beta_{a_{n}}}{a_{n}}$
$\qquad+131c\lambda\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$ (4.3)
where $w_{f}(s)=\left\{\sup\left|f(x)-f(y)\right|:x,y\in[0,1]^{d},\left\|x-y\right\|_{2}\leq s\right\}$ for any continuous function $f$ and $w_{\kappa}(s)=[w_{\kappa^{(1)}}(s)\ \cdots\ w_{\kappa^{(K)}}(s)]^{\top}$, $\beta_{a_{n}}$ is the $a_{n}$-th $\beta$-mixing coefficient of samples defined in Definition 5.3, $c$ is a positive constant and
$\lambda=\left\{\begin{array}[]{ll}1,&\ell_{1}\text{\ loss},\\ \frac{1}{\epsilon},&\text{CE loss},\\ -A,&\text{RCE loss}.\end{array}\right.$
Proof.
We defer the proof to Section 5. ∎
Remark 4.2.
In general, both the optimal classifier and the learned classifier within the framework of deep neural networks aim to maximize separability between different classes while minimizing in-class distances, which enables the classifier to achieve maximum margin classification. Furthermore, in $K$-class classification tasks, the theory of neural collapse for the CE loss [ZDZ+21, TKVB22, Liu24] suggests that the last-layer classifiers of well-trained deep neural networks, when trained on a balanced dataset, correspond to the $K$ vertices of the Simplex Equiangular Tight Frame (ETF). This framework ensures both concentration within classes and separation between different classes. Although we do not concentrate on $f_{0}:[0,1]^{d}\to\mathcal{E}=\{e_{1},\cdots,e_{K}\}$ in relation to the CE loss, the results on the empirical risk for the true distribution hold true by Theorem 4.1, provided that the outputs of the true classifier and the optimal classifier are sufficiently separated from the margins.
In the proof of the error bound for the excess empirical risk minimizer, we dissect it into two components: the statistical error and the approximation error. On the right-hand side of inequalities (4.2) and (4.1), the first two terms represent the upper bound of the statistical error, whereas the third term signifies the upper bound of the approximation error.
The second terms on the right-hand side of inequalities (4.2) and (4.1) arise due to the dependence among the data points. When the data points are independent, this dependency term diminishes, as the observations no longer influence each other. In such cases, the error bounds may simplify and potentially tighten, reflecting the reduced complexity due to the absence of interdependencies among the data.
In general, the inequality $|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq|\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})|$ holds because $\hat{f^{\eta}_{n}}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ is chosen as the empirical risk minimizer trained on a dataset with noisy labels. Meanwhile, $f_{0}$ denotes the true classifier.
Equation (4.2) in Theorem 4.1 is based on the noise-tolerant property of the loss function, which guarantees $\hat{f^{\eta}_{n}}=\hat{f_{n}}$. That is, $\mathcal{L}(\hat{f^{\eta}_{n}})-\mathcal{L}(f_{0})=\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})$. It is important to note that both the $\ell_{1}$ loss and the RCE loss exhibit noise-tolerant properties. However, for the CE loss, this noise-tolerant property holds only when $f$ is a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$. We will discuss this further in Subsection 5.1.
Additionally, Assumption 4.1 is required only for the CE and RCE losses; it is not needed for the $\ell_{1}$ loss at all.
Furthermore, we have
$w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\leq O((N^{2}M^{2}\log_{3}(N+2))^{-1/d})$
for any uniformly continuous functions $f(x)$.
Lastly, $\beta_{a_{n}}$ tends to zero as $a_{n}$ approaches infinity. Thus, $\beta_{a_{n}}$ can be made small by choosing $a_{n}$ sufficiently large.
5 Proof of Theorem 4.1
In this subsection, we provide a detailed proof of Theorem 4.1. In Subsection 5.1, we discuss the noise-tolerant properties of the loss function, demonstrating that the empirical minimizer on noisy data aligns with that on the true data. In Subsection 5.2, we estimate the error bounds of the excess risk for different loss functions, categorizing them into statistic error and approximation error. We analyze these two types of errors in Subsections 5.3 and 5.4.
5.1 Noise-tolerance
In the following, we will demonstrate that the empirical risk minimizer for the noisy data is the same as that for the true data, which is referred as noise-tolerant.
Definition 5.1.
A loss function $\ell(\cdot,\cdot)$ is referred to noise-tolerant if $\hat{f^{\eta}_{n}}=\hat{f_{n}}$, where $\hat{f^{\eta}_{n}},\hat{f_{n}}$ are the empirical risk minimizers for the noisy data and true data, respectively, as defined in (3.3).
To prove the noise-tolerant property, it is essential to define the symmetric loss function. We claim that a loss function is noise-tolerant if it is a symmetric loss function, which is stated in Theorem 5.1.
Definition 5.2 (Equation (2) in [GKS17]).
If a loss function $\ell(\cdot,\cdot)$ satisfies the following condition for a constant $C_{0}$, we refer to it as a symmetric loss function:
$\sum_{j=1}^{K}\ell(f(x),e_{j})=C_{0},\quad\forall f\in{\mathcal{F}}.$ (5.1)
Theorem 5.1.
Let $K$ be the number of classes in the classification problem with noisy labels. We assume that the noisy labels are derived from the true labels through symmetric noise, as defined in (4.1), with the noise rate $\eta<\frac{K-1}{K}$. When a loss function satisfies (5.1), it is noise-tolerant.
Proof.
Let $\{(x_{i},y_{i},y^{\eta}_{i})\}^{n}_{i=1}$ be $n$ random samples from the distribution of $(X,Y,Y^{\eta})$. For any $f\in{\mathcal{F}}$,
$\mathcal{L}^{\eta}_{n}(f)$ $=\frac{1}{n}\sum^{n}_{i=1}\ell(f(x_{i}),y^{\eta}_{i})=\frac{1}{n}\sum^{n}_{i=1}\left[(1-\eta)\ell(f(x_{i}),y_{i})+\frac{\eta}{K-1}\sum^{K}_{j=1}\sum_{y_{i}\neq e_{j}}\ell(f(x_{i}),e_{j})\right]$
$=(1-\eta)\mathcal{L}_{n}(f)+\frac{\eta}{K-1}(C_{0}-\mathcal{L}_{n}(f))=\frac{C_{0}\eta}{K-1}+\left(1-\frac{\eta K}{K-1}\right)\mathcal{L}_{n}(f),$
where the third equality is from Equation (5.1). Since $\eta<\frac{K-1}{K}$ and $\hat{f_{n}}$ is the empirical minimizer for the true data in (3.3), then we have
$\mathcal{L}^{\eta}_{n}(\hat{f}_{n})-\mathcal{L}^{\eta}_{n}(f)=\left(1-\frac{\eta K}{K-1}\right)(\mathcal{L}_{n}(\hat{f}_{n})-\mathcal{L}_{n}(f))\leq 0,\ \forall f\in{\mathcal{F}},$
which implies that $\hat{f}_{n}=\hat{f}^{\eta}_{n}$. That is, the loss function $\ell(\cdot,\cdot)$ is noise-tolerant. ∎
For the loss functions
$l(\hat{y},y)=\left\{\begin{array}[]{ll}\|\hat{y}-y\|_{1},&\ell_{1}\text{ loss},\\ -\sum_{i=1}^{k}y_{i}\log(\hat{y_{i}}),&\text{CE loss},\\ -\sum_{i=1}^{k}\hat{y_{i}}\log(y_{i}),&\text{RCE loss},\end{array}\right.$
we have the following remarks:
Remark 5.2.
Both the $\ell_{1}$ loss and the RCE loss are symmetric loss functions when we choose $f$ as a function from ${\mathbb{R}}^{d}$ to $\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0,\sum^{K}_{i=1}=1\}$ [GKS17, WMC+19]. For CE loss, it is also a symmetric loss function when we set $f$ to be a function mapping ${\mathbb{R}}^{d}$ to =$\mathcal{E}=\{e_{1},\cdots,e_{K}\}$.
5.2 Error analysis
In this subsection, we analyze the error bounds for different types of loss functions, including the $\ell_{1}$ loss, the CE and RCE losses. We start from the Lipschitz-type property for the CE and RCE losses.
Lemma 5.1.
When Assumption 4.1 holds, both the CE loss function and the RCE loss function satisfy the following property
$\left|\ell(a,q)-\ell(b,q)\right|\leq\lambda\|a-b\|_{1},\ \forall a,b,q\in\mathcal{Y},$
where
$\lambda=\left\{\begin{array}[]{ll}\frac{1}{\epsilon},&\mathrm{CE\ loss},\\ -A,&\mathrm{RCE\ loss}.\end{array}\right.$ (5.2)
Proof.
For CE loss function $\ell(a,q)=-\sum_{i=1}^{K}q_{i}\log a_{i}$, $\ell(b,q)=-\sum_{i=1}^{K}q_{i}\log b_{i}$, we have
$\left|\ell(a,q)-\ell(b,q)\right|$ $=\left|\sum_{i=1}^{K}q_{i}\log a_{i}-\sum_{i=1}^{K}q_{i}\log b_{i}\right|=\left|\sum_{i=1}^{K}q_{i}[\log a_{i}-\log b_{i}]\right|$
$\leq\sum_{i=1}^{K}\left|q_{i}\right|\left|\log a_{i}-\log b_{i}\right|=\sum_{i=1}^{K}\left|q_{i}\right|(\log x)^{\prime}|_{x=\xi_{i}}\left|a_{i}-b_{i}\right|,(\exists\xi_{i}\in[a_{i},b_{i}])$
$\leq\frac{1}{\epsilon}\|q\|_{2}\|a-b\|_{2}\leq\frac{1}{\epsilon}\|a-b\|_{1}.$
where the third equality is from the Lagrange mean value theorem and $\max_{x}(\log x)^{\prime}=1/\epsilon$ when Assumption 4.1 holds.
For RCE loss function $\ell(a,q)=-\sum_{i=1}^{K}a_{i}\log q_{i}$, $\ell(b,q)=-\sum_{i=1}^{K}b_{i}\log q_{i}$, we directly calculate :
$\left|\ell(a,q)-\ell(b,q)\right|$ $=\left|\sum_{i=1}^{K}a_{i}\log q_{i}-\sum_{i=1}^{K}b_{i}\log q_{i}\right|=\left|\sum_{i=1}^{K}\log q_{i}(a_{i}-b_{i})\right|$
$\leq\sum_{i=1}^{K}\left|\log q_{i}\right|\left|(a_{i}-b_{i})\right|\leq\max_{i}|\log q_{i}|\|a-b\|_{1}\leq-A\|a-b\|_{1},$
where we use the fact that $\max_{i}|\log q_{i}|\leq-A$ based on Assumption 4.1. ∎
Lemma 5.2.
Let $f_{0}:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0,\sum^{K}_{i=1}=1\}$ represent the true classifier, and $\hat{f_{n}}$ denote the empirical risk minimizer on the true data. Let $S=\{(x_{1},y_{1}),\cdots,(x_{n},y_{n})\in X\times Y\}$. We have the following
(1) When the loss function is the $\ell_{1}$ loss
$|\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})|\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}.$ (5.3)
(2) When Assumption 4.1 holds and the loss function is either the CE loss or the RCE loss
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$ (5.4)
where $\lambda$ is defined in (5.2).
Proof.
Let $\tilde{f}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ satisfying $\|\tilde{f}-f_{0}\|_{L^{1}([0,1]^{d})}\leq\|f-f_{0}\|_{L^{1}([0,1]^{d})},\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$, then
$|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|$ $=\mathbb{E}_{S}(|\mathcal{L}(\hat{f}_{n})-\mathcal{L}(f_{0})|)$ (5.5)
$=\mathbb{E}_{S}(|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})+\mathcal{L}_{n}(\hat{f}_{n})-\mathcal{L}_{n}(\tilde{f})+\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})+\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|)$
$\leq\mathbb{E}_{S}\left|\mathcal{L}(\hat{f}_{n})-\mathcal{L}_{n}(\hat{f}_{n})\right|+\mathbb{E}_{S}\left|\mathcal{L}_{n}(\tilde{f})-\mathcal{L}(\tilde{f})\right|+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|$
$\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\left|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})\right|.$
Here, we utilize the fact that $\mathcal{L}_{n}(\hat{f}_{n})\leq\mathcal{L}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ in the first inequality of (5.5). For the $\ell_{1}$ loss function, we have
$|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$ $=|\mathbb{E}_{X,Y}\|\tilde{f}(X)-Y\|_{1}-\mathbb{E}_{X,Y}\|{f_{0}}(X)-Y\|_{1}|\leq\mathbb{E}_{X,Y}[|\|\tilde{f}(X)-Y\|_{1}-\|{f_{0}}(X)-Y\|_{1}|]$ (5.6)
$\leq\mathbb{E}_{X}\mathbb{E}_{Y|X}[\|\tilde{f}(X)-f_{0}(X)\|_{1}]=\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}=\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}$
$=\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from the triangle inequality. For the CE loss function or the RCE loss function, we have
$|\mathcal{L}(\tilde{f})-\mathcal{L}(f_{0})|$ $=\left|\mathbb{E}_{X,Y}\left[\ell(\tilde{f}(X),Y)\right]-\mathbb{E}_{X,Y}\left[\ell(f_{0}(X),Y)\right]\right|\leq\lambda\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}$ (5.7)
$\leq\lambda\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}=\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from Lemma 5.1 and $\lambda$ is defined in (5.2). Combining (5.5), (5.6) and (5.7), we conclude the result. ∎
For the excess risk about the empirical risk minimizer on the noisy distribution $\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})$, we have the following similar results.
Lemma 5.3.
Let $f_{0}$ represent the true classifier, and $\hat{f^{\eta}_{n}}$ denote the empirical risk minimizer on the noisy data. Let $S^{\eta}=\{(x_{1},y^{\eta}_{1}),\cdots,(x_{n},y^{\eta}_{n})\in X\times Y^{\eta}\}$. We have the following
(1) When the loss function is the $\ell_{1}$ loss
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}.$ (5.8)
(2) When Assumption 4.1 holds and the loss function is either the CE los or the RCE loss
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$ (5.9)
where $\lambda$ is defined in (5.2).
Proof.
Let $\tilde{f}\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ satisfying $\|\tilde{f}-f_{0}\|_{L^{1}([0,1]^{d})}\leq\|f-f_{0}\|_{L^{1}([0,1]^{d})},\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$, then
$|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|$ $=\mathbb{E}_{S^{\eta}}(|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}(f_{0})|)$ (5.10)
$=\mathbb{E}_{S^{\eta}}(|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})+\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\tilde{f})+\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})+\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|)$
$\leq\mathbb{E}_{S^{\eta}}\left|\mathcal{L}^{\eta}(\hat{f^{\eta}_{n}})-\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\right|+\mathbb{E}_{S}\left|\mathcal{L}^{\eta}_{n}(\tilde{f})-\mathcal{L}^{\eta}(\tilde{f})\right|+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|$
$\leq 2\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]+\left|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})\right|.$
Here, we utilize the fact that $\mathcal{L}^{\eta}_{n}(\hat{f^{\eta}_{n}})\leq\mathcal{L}^{\eta}_{n}(f),\ \forall f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})$ in the first inequality of (5.10). For the $\ell_{1}$ loss function, we have
$|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$ $=|\mathbb{E}_{X,Y^{\eta}}\|\tilde{f}(X)-Y^{\eta}\|_{1}-\mathbb{E}_{X,Y^{\eta}}\|{f_{0}}(X)-Y^{\eta}\|_{1}|$ (5.11)
$\leq\mathbb{E}_{X,Y^{\eta}}[|\|\tilde{f}(X)-Y^{\eta}\|_{1}-\|{f_{0}}(X)-Y^{\eta}\|_{1}|]$
$\leq\mathbb{E}_{X}\mathbb{E}_{Y^{\eta}|X}[\|\tilde{f}(X)-f_{0}(X)\|_{1}]=\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}=\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}$
$=\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from the triangle inequality. For the CE loss function or the RCE loss function, we have
$|\mathcal{L}^{\eta}(\tilde{f})-\mathcal{L}^{\eta}(f_{0})|$ $=\mathbb{E}_{X,Y^{\eta}}\left[\ell(\tilde{f}(X),Y^{\eta})\right]-\mathbb{E}_{X,Y^{\eta}}\left[\ell(f_{0}(X),Y^{\eta})\right]\leq\lambda\mathbb{E}_{X}\|\tilde{f}(X)-f_{0}(X)\|_{1}$ (5.12)
$\leq\lambda\left\|\tilde{f}-f_{0}\right\|_{L^{1}([0,1]^{d})}=\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})},$
where the second inequality is from Lemma 5.1 and $\lambda$ is defined in (5.2). Combining (5.10), (5.11) and (5.12), we conclude the result. ∎
We denote $\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]$ or $\mathbb{E}_{S^{\eta}}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]$ as the statistical error and $\left\|f-f_{0}\right\|_{L^{1}([0,1]^{d})}$ as the approximation error. In the following subsections, we will explore both types of errors.
5.3 Statistical error
In practice, the samples in the dataset may not be independent. In this subsection, we focus on a dependent (mixing) sequence. To bound the statistical errors, we need several definitions and lemmas below.
Definition 5.3 (Definition from Subsection 2.2 in [Yu94]).
Let $\left\{x_{t}\right\}_{t\geq 1}$ be a stochastic process, and define $x^{i:j}=(x_{i},\dots,x_{j})$ for $i\leq j$, where $j$ can be infinite. Let $\sigma(x^{i:j})$ denote the $\sigma$-algebra generated by $x^{i:j}$ for $i\leq j$. The $s$-th $\beta$-mixing coefficient of $\{x_{t}\}_{t\geq 1}$, denoted by $\beta_{s}$, is defined as
$\beta_{s}=\sup_{t\geq 1}\mathbb{E}\left[\sup_{B\in\sigma(x^{t+s:\infty})}\left|\mathrm{Prob}(B|\sigma(x^{1:t}))-\mathrm{Prob}(B)\right|\right].$ (5.13)
We say that $\{x_{t}\}_{t\geq 1}$ is $\beta$-mixing if $\beta_{s}\to 0$ as $s\to\infty$.
Definition 5.4 (Independent Block Sequence in [Yu94]).
Let $(a_{n},\mu_{n})$ be an integer pair such that $n=a_{n}\mu_{n}$. We divide the strictly stationary sequence $\{\hat{x}_{i}\}_{i=1}^{2n}$ into $2\mu_{n}$ blocks, each of length $a_{n}$. For $1\leq j\leq\mu_{n}$, define
$G_{j}:=\left\{i:2(j-1)a_{n}+1\leq i\leq(2j-1)a_{n}\right\},\ \hat{x}_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}_{i}.$
Let $\hat{x}_{i}^{\prime}$ be an i.i.d. copy of $\hat{x}_{i}$ for $i=1,\cdots,2n$. That is,
$\Xi_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}^{\prime}_{i}$
is a sequence of identically distributed independent blocks such that the sequence is independent of $\{\hat{x}_{i}\}^{2n}_{i=1}$ and each block has the same distribution as a block from the original sequence. We call $\Xi_{1,a_{n}}$ the independent block $a_{n}$-sequence (IB sequence).
Lemma 5.4 (Lemma from Subsection 4.1 in [Yu94]).
Let $x_{1,a_{n}}$ and $\Xi_{1,a_{n}}$ be the mixing and independent block sequences defined in Definition 5.4, whose distributions are $\mathcal{T}$ and $\tilde{\mathcal{T}}$, respectively. For any measurable function $g$ on ${\mathbb{R}}^{\mu_{n}a_{n}}$ with bound $\tilde{M}$,
$\left|\mathcal{T}g(x_{1,a_{n}})-\tilde{\mathcal{T}}g(\Xi_{1,a_{n}})\right|\leq\tilde{M}(\mu_{n}-1)\beta_{a_{n}},$
where $(a_{n},\mu_{n})$ is an integer pair with $n=2a_{n}\mu_{n}$, and $\beta_{a_{n}}$ is defined in (5.13).
Lemma 5.5 (Lemma 5 in [MZ03]).
Let $\left\{g_{i}(\theta)\right\}$ and $\left\{h_{i}(\theta)\right\}$ be sets of functions defined for all $\theta$ in some domain $\Theta$. If $\left|g_{i}(\theta)-g_{i}(\theta^{\prime})\right|\leq\left|h_{i}(\theta)-h_{i}(\theta^{\prime})\right|$ holds for all $i,\theta,\theta^{\prime}$, then for any function $c(x,\theta)$ and probability distribution of x is over $X$, we have
$\mathbb{E}_{\theta}\mathbb{E}_{X}\sup_{\theta\in\Theta}\left\{c(X,\theta)+\sum_{i=1}^{n}\sigma_{i}g_{i}(\theta)\right\}\leq\mathbb{E}_{\theta}\mathbb{E}_{X}\sup_{\theta\in\Theta}\left\{c(X,\theta)+\sum_{i=1}^{n}\sigma_{i}h_{i}(\theta)\right\}.$
Lemma 5.6 (Theorem 2 in [GRS18]).
Let ${\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}})$ represent a class of neural networks, characterized by an input dimension $d$, an output dimension of $1$, and a width and depth denoted by ${\mathcal{W}}$ and ${\mathcal{D}}$, respectively. Define $W_{j}$ as the weight matrix of the $j$-th layer, satisfying $\|W_{j}\|_{1,\infty}\leq M(j)$ for all $j=1,\cdots,{\mathcal{D}}+1$, where $\|W_{j}\|_{1,\infty}$ denotes the infinity norm of the 1-norms of the columns of $W_{j}$. Let $\hat{S}=\{x_{i}\}^{n}_{i=1}$ represent $n$ input features such that $\|x_{i}\|_{2}\leq B,\forall i$ and $\sigma_{i},i=1,\cdots,n$ are i.i.d. Rademacher random variables that are independent of $\hat{S}$. Then
$\begin{split}\mathbb{E}_{\{x_{i},\sigma_{i}\}_{i=1}^{n}}\left[\sup_{f\in\mathcal{F}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}f(x_{i})\right|\right]\leq&\dfrac{2}{n}\prod\limits_{j=1}^{{\mathcal{D}}+1}M(j)\cdot\sqrt{{\mathcal{D}}+2+\log d}\cdot\sqrt{\max_{i\in\{1,\cdots,n\}}\|x_{i}\|^{2}_{2}}\\ \leq&\frac{2B\sqrt{{\mathcal{D}}+2+\log d}\cdot\prod_{j=1}^{{\mathcal{D}}+1}M(j)}{\sqrt{n}}.\end{split}$
Theorem 5.3.
Let ${\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ represent a class of the norm-constrained neural network, characterized by an input dimension $d$, an output dimension of $K$, a width and depth denoted by ${\mathcal{W}}$ and ${\mathcal{D}}$, and a constraint ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\left\{\|(W_{i},b_{i})\|_{1,\infty},1\right\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer for $i=0,1,\cdots,{\mathcal{D}}$. Let $S=\left\{z_{i}=(x_{i},y_{i})\right\}_{i=1}^{n}$ be a strictly stationary $\beta$-mixing sequence and $(a_{n},\mu_{n})$ be an integer pair such that $n=a_{n}\mu_{n}$. Suppose the loss function we consider satisfying $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$, where $\lambda>0$ is a given constant. Then we have
$\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]\leq\frac{4\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{2\lambda Kn\beta_{a_{n}}}{a_{n}}.$
Proof.
Let $\tilde{z_{i}}$ denote an independent copy of $z_{i}$, and denote $\tilde{S}=\{\tilde{z}_{i}\}^{n}_{i=1}$. Set $z_{i}=(x_{i},y_{i})$, $\tilde{z}_{i}=(\tilde{x}_{i},\tilde{y}_{i})$ and define $\ell_{f}(z_{i})=\ell(f(x_{i}),y_{i})$ and $\ell_{f}(\tilde{z}_{i})=\ell(f(\tilde{x}_{i}),\tilde{y}_{i})$. Then we have
$\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,k}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]=\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(z_{i})-\mathbb{E}_{\tilde{S}}[\ell_{f}(\tilde{z_{i}})]\right|\right]$
$=$ $\mathbb{E}_{S}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{S}}[\ell_{f}(z_{i})]-\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{S}}[\ell_{f}(\tilde{z_{i}})]\right|\right]$
$\leq$ $\mathbb{E}_{S,\tilde{S}}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(z_{i})-\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(\tilde{z_{i}})\right|\right]$
$=$ $2\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}\ell_{f}(\hat{z}_{i})\right|\right]$
$\leq$ $2\lambda K\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in\mathcal{F}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}f(\hat{x}_{i})\right|\right],$ (5.14)
where $\hat{S}=\{\hat{z}_{1},\cdots,\hat{z}_{2n}\}=S\cup\tilde{S}$, $\sigma_{i},i=1,\cdots,2n$ are i.i.d. Rademacher random variables that are independent of $\hat{S}$ and $\sigma=\{\sigma_{i}\}^{2n}_{i=1}$. For the last inequality, we combine Lemma 5.5 and the fact that $\ell_{f}(\hat{z}_{i})-\ell_{f}(\hat{z}^{\prime}_{i})\leq\lambda\|f(\hat{x}_{i})-f(\hat{x}^{\prime}_{i})\|_{1}\leq K\lambda|f_{j}(\hat{x}_{i})-f_{j}(\hat{x}^{\prime}_{i})|$ with $j=\mathop{\arg\max}_{j\in\{1,\cdots,K\}}|f_{j}(\hat{x}_{i})-f_{j}(\hat{x}^{\prime}_{i})|$.
We partition the strictly stationary sequence $\{\hat{x}_{i}\}_{i=1}^{2n}$ into $2\mu_{n}$ blocks, each of length $a_{n}$. For $1\leq j\leq\mu_{n}$, define
$G_{j}:=\left\{i:2(j-1)a_{n}+1\leq i\leq(2j-1)a_{n}\right\},\ H_{j}:=\left\{i:(2j-1)a_{n}+1\leq i\leq 2ja_{n}\right\}.$
In addition, we define
$\hat{x}_{1,a_{n}}$ $=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}_{i},\ \sigma_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\sigma_{i},\ y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})=\sum_{i\in G_{j}}\sigma_{i}f(\hat{x}_{i}),$
$\hat{x}_{2,a_{n}}$ $=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in H_{j}}\hat{x}_{i},\ \sigma_{2,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in H_{j}}\sigma_{i},\ y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})=\sum_{i\in H_{j}}\sigma_{i}f(\hat{x}_{i}).$
Let $\hat{x}_{i}^{\prime}$ be an i.i.d. copy of $\hat{x}_{i}$ for $i=1,\cdots,2n$. Define
$\Xi_{1,a_{n}}=\bigcup^{\mu_{n}}_{j=1}\bigcup_{i\in G_{j}}\hat{x}^{\prime}_{i},\ W_{j,g}(\Xi_{1,a_{n}},\sigma_{1,a_{n}})=\sum_{i\in G_{j}}\sigma_{i}f(\hat{x}^{\prime}_{i}).$
Then $\Xi_{1,a_{n}}$ is the IB sequence of $x_{1,a_{n}}$. We have
$\mathbb{E}_{\hat{S},\sigma}\left[\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{i=1}^{2n}\sigma_{i}f(\hat{x}_{i})\right|\right]$
$=$ $\mathbb{E}_{\hat{S},\sigma}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})+\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})\right|\right)$
$\leq$ $\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})\right|\right)$
$\qquad+\mathbb{E}_{\hat{x}_{2,a_{n}},\sigma_{2,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{2,a_{n}},\sigma_{2,a_{n}})\right|\right)$
$=$ $2\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{2n}\sum_{j=1}^{\mu_{n}}y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})\right|\right)$
$=$ $\mathbb{E}_{\hat{x}_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{\mu_{n}}\sum_{j=1}^{\mu_{n}}\frac{y_{j,f}(\hat{x}_{1,a_{n}},\sigma_{1,a_{n}})}{a_{n}}\right|\right)$
$\leq$ $\mathbb{E}_{\Xi_{1,a_{n}},\sigma_{1,a_{n}}}\left(\sup_{f\in{\mathcal{F}}_{d,1}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\frac{1}{\mu_{n}}\sum_{j=1}^{\mu_{n}}\frac{W_{j,f}(\Xi_{1,a_{n}},\sigma_{1,a_{n}})}{a_{n}}\right|\right)+\mu_{n}\beta_{a_{n}}$
$\leq$ $\frac{2{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{a_{n}\sqrt{\mu_{n}}}+\mu_{n}\beta_{a_{n}}=\frac{2{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{n\beta_{a_{n}}}{a_{n}},$ (5.15)
where the second inequality to last is from Lemma 5.4, the last inequality is from Lemma 5.6 and the last equality is from $n=a_{n}\mu_{n}$. Combining (5.3) and (5.3), we conclude the result. ∎
Remark 5.4.
The condition $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$ holds for the $\ell_{1}$ loss by the triangle inequality. For the CE or the RCE loss, this condition holds when Assumption 4.1 is satisfied, as derived from Lemma 5.1. The values of $\lambda$ for different loss functions are
$\lambda=\left\{\begin{array}[]{ll}1,&\ell_{1}\text{\ loss},\\ \frac{1}{\epsilon},&\text{CE loss},\\ -A,&\text{RCE loss}.\end{array}\right.$
Theorem 5.5.
Let ${\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})$ represent a class of the norm-constrained neural network, characterized by an input dimension $d$, an output dimension of $K$, a width and depth denoted by ${\mathcal{W}}$ and ${\mathcal{D}}$, and a constraint ${\mathcal{M}}>0$ satisfying $\|(W_{{\mathcal{D}}},b_{{\mathcal{D}}})\|_{1,\infty}\prod_{i=0}^{{\mathcal{D}}-1}\max\left\{\|(W_{i},b_{i})\|_{1,\infty},1\right\}\leq{\mathcal{M}}$, where $W_{i}\in\mathbb{R}^{p_{i+1}\times p_{i}}$ is the weight matrix, and $b_{i}\in\mathbb{R}^{p_{i+1}}$ is the bias vector in the $i$-th layer for $i=0,1,\cdots,{\mathcal{D}}$. We consider the noisy dataset $S^{\eta}$ as a strictly stationary $\beta$-mixing sequence with the sample size $n$ satisfying $n=a_{n}\mu_{n}$, where $(a_{n},\mu_{n})$ is an integer pair. Suppose the loss function we consider satisfying $|\ell(a,q)-\ell(b,q))|\leq\lambda\|a-b\|_{1}$, where $\lambda>0$ is a given constant. Then we have
$\mathbb{E}_{S^{\eta}}\left[\sup_{f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}},{\mathcal{M}})}\left|\mathcal{L}^{\eta}(f)-\mathcal{L}^{\eta}_{n}(f)\right|\right]\leq\frac{4\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{\sqrt{na_{n}}}+\frac{2\lambda Kn\beta_{a_{n}}}{a_{n}}.$
Proof.
The proof is very similar to that in Theorem 5.3, so we will omit it to save space. ∎
5.4 Approximate error
First of all, we give a definition and some lemmas, which will be useful for the proof of approximate error bound. To estimate the boundary of the approximation error, we utilize certain proofs from [SYZ22] for functions within $\mathcal{F}_{d,1}({\mathcal{W}},{\mathcal{D}})$ and extend these proofs to encompass the neural network class $\mathcal{F}_{d,K}({\mathcal{W}},{\mathcal{D}})$.
Definition 5.5.
Let $f=[f^{(1)}\ \cdots\ f^{(K)}]^{\top}$ be a map from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$, where $f^{(i)}\in C([0,1]^{d})$. Define $w_{f}(s)$ as the continuous module,
$w_{f}(s):=\left[w_{f^{(1)}}(s)\ \cdots\ w_{f^{(K)}}(s)\right]^{\top},\ \forall s\geq 0,$ (5.16)
where $w_{f^{(i)}}(s)=\left\{\sup\left|f^{(i)}(x)-f^{(i)}(y)\right|:x,y\in[0,1]^{d},\left\|x-y\right\|_{2}\leq s\right\}$. Especially, we have
$w_{f}(s\cdot j)\leq j\cdot w_{f}(s),\ \forall j\in\mathbb{N}^{+},\forall s\geq 0.$ (5.17)
Lemma 5.7 (Lemma 25 in [FJK+23]).
Let $\mathcal{F}_{d,K}({\mathcal{W}},{\mathcal{D}})$ denote the class of functions generated by a ReLU neural network with input dimension $d$, output dimension $K$, width ${\mathcal{W}}$ and depth ${\mathcal{D}}$. Suppose $g_{1}\in\mathcal{F}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1})$, $g_{2}\in\mathcal{F}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2})$, then the following properties hold:
(1) If $d_{1}=d_{2}$, $K_{1}=K_{2}$, ${\mathcal{W}}_{1}\leq{\mathcal{W}}_{2}$, ${\mathcal{D}}_{1}\leq{\mathcal{D}}_{2}$, then ${{\mathcal{F}}}_{d_{1},K_{1}}({\mathcal{W}}_{1},{\mathcal{D}}_{1})\subseteq{{\mathcal{F}}}_{d_{2},K_{2}}({\mathcal{W}}_{2},{\mathcal{D}}_{2})$.
(2) If $K_{1}=d_{2}$, then $g_{2}\circ g_{1}\in\mathcal{F}_{d_{1},K_{2}}(\max\left\{{\mathcal{W}}_{1},{\mathcal{W}}_{2}\right\},{\mathcal{D}}_{1}+{\mathcal{D}}_{2})$.
Lemma 5.8.
Under the same assumption of Lemma 5.7, define $x=(x_{1},x_{2})^{\top}$ and $g(x)=(g_{1}(x_{1}),g_{2}(x_{2}))$, then $g\in\mathcal{F}_{d_{1}+d_{2},K_{1}+K_{2}}({\mathcal{W}}_{1}+{\mathcal{W}}_{2},\max\left\{{\mathcal{D}}_{1},{\mathcal{D}}_{2}\right\})$.
Lemma 5.9 (Proposition 3.1 in [SYZ22]).
For any $N,M,d\in\mathbb{N}^{+}$, define $J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor$ and $\delta\in(0,\frac{1}{3J}]$. Then there exists a function $\psi\in\mathcal{F}_{1,1}\left(8\left\lfloor N^{1/d}\right\rfloor+3,2\left\lfloor M^{1/d}\right\rfloor+5\right)$, such that
$\psi(x)=j,\quad x\in\left[\frac{j}{J},\frac{j+1}{J}-\delta\cdot\mathds{1}_{\left\{j\leq J-2\right\}}\right],\quad j=0,1,\cdots,J-1.$
Lemma 5.10 (Proposition 3.2 in [SYZ22]).
For any $\epsilon>0$ and any $N,M\in\mathbb{N}^{+}$, define $J$ to be a positive integer satisfying $J\leq N^{2}M^{2}\left\lfloor\log_{3}(N+2)\right\rfloor$. Assume $f(x_{j})>0$ for $j=0,1,\cdots,J-1$ are samples with
$\left|f(x_{j})-f(x_{j-1})\right|\leq\epsilon,\ j\in[J].$
Then there exists $\phi\in\mathcal{F}_{1,1}(16N+30,6M+10)$ such that
[…]
Theorem 5.6.
For any continuous map $f:[0,1]^{d}\to{\mathbb{R}}^{K}$ and any $N,M\in\mathbb{N}^{+}$, there exists a function
$\phi\in\mathcal{F}_{d,K}\left(\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\},11M+18\right),$
such that $\left\|\phi\right\|_{L^{\infty}({\mathbb{R}}^{d})}\leq\left|f(\mathbf{0})\right|+w_{f}(\sqrt{d})$ and
$\left|\phi(x)-f(x)\right|\leq\left\{\begin{array}[]{ll}130\sqrt{d}w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right),&\forall x\in[0,1]^{d}\backslash\Omega([0,1]^{d},J,\delta),\\ 2w_{f}(\sqrt{d}),&\forall x\in\Omega([0,1]^{d},J,\delta),\end{array}\right.$
where $J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor$, $\forall\delta\in(0,\frac{1}{3J}]$ and
$\Omega([0,1]^{d},J,\delta)=\bigcup_{i=1}^{d}\left\{x=[x_{1},x_{2},\cdots,x_{d}]^{\top}:x_{i}\in\bigcup_{j=1}^{J}(j/J-\delta,j/J)\right\}.$
Proof.
(1) We divide $[0,1]^{d}$ into two parts. Let $p=[p_{1},p_{2},\cdots,p_{d}]^{\top}\in[J]^{d}=\{0,1,\cdots,J-1\}^{d}$, define $x_{p}=p/J$ and
$Q_{p}=\left\{x=[x_{1},x_{2},\cdots,x_{d}]^{\top}:x_{i}\in\left[\frac{p_{i}}{J},\frac{p_{i}+1}{J}-\delta\cdot\mathds{1}_{\left\{p_{i}\leq J-2\right\}}\right],i=1,2,\cdots,d\right\}.$
Then $[0,1]^{d}$ can be represented as
$[0,1]^{d}=\cup_{p\in[J]^{d}}Q_{p}\bigcup\Omega([0,1]^{d},J,\delta).$
(2) We build a mapping $\phi_{1}$ from $Q_{p}$ to $p$. There exists a function $\psi\in{\mathcal{F}}_{1,1}\left(8\left\lfloor N^{1/d}\right\rfloor+3,2\left\lfloor M^{1/d}\right\rfloor+5\right)$ from Lemma 5.9, such that
$\psi(x_{j})=p_{j},\quad x_{j}\in\left[\frac{p_{j}}{J},\frac{p_{j}+1}{J}-\delta\cdot\mathds{1}_{\left\{p_{j}\leq J-2\right\}}\right],\quad p_{j}\in[J],\quad j=1,2,\cdots,d,$
where
$J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor.$ (5.18)
Then we define
$\phi_{1}(x):=[\psi(x_{1}),\psi(x_{2}),\cdots,\psi(x_{d})]^{\top}=p,\quad\forall x=[x_{1},x_{2},\cdots,x_{d}]^{\top}\in Q_{p}.$ (5.19)
(3) Define two auxiliary sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Construct a linear function $\psi_{1}$ on ${\mathbb{R}}^{d}$, which maps $[J]^{d}$ to $\mathcal{S}_{1}$ exactly. Let
$\psi_{1}(x):=\frac{x_{d}}{2J^{d}}+\sum_{i=1}^{d-1}\frac{x_{i}}{J^{i}},\ \forall x=[x_{1},x_{2},\cdots,x_{d}]\in\mathbb{R}^{d}$
be a linear mapping. Define
$\mathcal{S}_{1}:=\left\{\frac{i}{J^{d-1}}+\frac{j}{2J^{d}}:i=0,1,\cdots,J^{d-1}-1,j\in[J]\right\},$
$\mathcal{S}_{2}:=\left\{\frac{i}{J^{d-1}}+\frac{j+J}{2J^{d}}:i=0,1,\cdots,J^{d-1}-1,j\in[J]\right\}.$
It is easy to get
$\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\left\{1\right\}=\left\{\frac{j}{2J^{d}}:j=0,1,\cdots,2J^{d}\right\}$
and
$\bigcup_{p\in[J]^{d}}\psi_{1}(p)$ =$=\left\{\frac{p_{d}}{2J^{d}}+\sum_{i=1}^{d-1}\frac{p_{i}}{J^{i}}:p\in[J]^{d}\right\}=\left\{\frac{i}{J^{d-1}}+\frac{j}{2J^{d}}:i=0,1,\cdots,J^{d-1}-1,j\in[J]\right\}=\mathcal{S}_{1}.$
(4) We normalize $f$ as $\bar{f}$ and construct a function $h$ on $\mathcal{S}_{1}\cup\mathcal{S}_{2}\cup\{1\}$ such that it fulfills $h\circ\psi_{1}(p)=\bar{f}(x_{p})$. Since $\left\|x-\mathbf{0}\right\|_{2}=\sqrt{(x_{1}-0)^{2}+\cdots+(x_{d}-0)^{2}}\leq\sqrt{1+\cdots+1}=\sqrt{d}$, we have $\left|f(x)-f(\mathbf{0})\right|\leq w_{f}(\sqrt{d}),\ \forall x\in[0,1]^{d}$. Define $\bar{f}$ as $\bar{f}(x):=f(x)-f(\mathbf{0})+w_{f}(\sqrt{d}),\ \forall x\in[0,1]^{d}$, we have
$0\leq\bar{f}(x)\leq 2w_{f}(\sqrt{d}).$ (5.20)
Let $h$ be a continuous piecewise linear function with a set of breakpoints, the breakpoint being $\mathcal{S}_{1}\cup\left\{1\right\}$. Define
$h(\psi_{1}(p))=\bar{f}(x_{p}),\quad\forall p\in[J]^{d},\ h(1)=\bar{f}(\mathbf{1}),$
and let $h$ to be a linear function on each interval in $\{[\frac{i}{J^{d-1}}-\frac{J+1}{2J^{d}},\frac{i}{J^{d-1}}]:j\in[J],i=1,2,\cdots,J^{d-1}\}\bigcup\{[\frac{i}{J^{d-1}}+\frac{j}{2J^{d}},\frac{i}{J^{d-1}}+\frac{j+1}{2J^{d}}]:j\in[J],i=1,2,\cdots,J^{d-1}\}$, where $\mathbf{1}=[1,1,\cdots,1]^{\top}\in{\mathbb{R}}^{d}$. Note that $\{\{\frac{i}{J^{d-1}}-\frac{J+1}{2J^{d}},\frac{i}{J^{d-1}}\},i=1,2,\cdots,J^{d-1}\}\subseteq\mathcal{S}_{1}\cup\left\{1\right\}$ and
$\mathcal{S}_{2}\subseteq{\textstyle\bigcup_{i=1}^{J^{d-1}}}\left[\frac{i}{J^{d-1}}-\frac{J+1}{2J^{d}},\frac{i}{J^{d-1}}\right],$
then $h$ satisfies
$\left|h\left(\frac{j}{2J^{d}}\right)-h\left(\frac{j-1}{2J^{d}}\right)\right|\leq\max\left\{w_{f}(1/J),w_{f}(\sqrt{d})/J\right\}\leq w_{f}\left(\frac{\sqrt{d}}{J}\right),\quad j=1,2,\cdots,2J^{d},$
where the last inequality is from (5.17). In addition, we have from (5.20) and the construction of $h$
$0\leq h\left(\frac{j}{2J^{d}}\right)\leq 2w_{f}(\sqrt{d}),\quad j=0,1,\cdots,2J^{d}.$
(5) Construct a function $\psi_{2}$ to approximate the function $h$. From (5.18), we have
$2J^{d}=2\left(\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor\right)^{d}\leq 2N^{2}M^{2}\left\lfloor\log_{3}(N+2)\right\rfloor\leq N^{2}\left\lceil\sqrt{2}M\right\rceil^{2}\left\lfloor\log_{3}(N+2)\right\rfloor,$
then there exists a function $\tilde{\psi}_{2}\in\mathcal{F}_{1,K}\left(16NK+30K,6\left\lceil\sqrt{2}M\right\rceil+10\right)$ from Lemma 5.10, such that
$\left|\tilde{\psi}_{2}(j)-h\left(\frac{j}{2J^{d}}\right)\right|\leq w_{f}\left(\frac{\sqrt{d}}{J}\right),\quad j=0,1,\cdots,2J^{d}-1,$
and
$0\leq\tilde{\psi}_{2}(\alpha)\leq\max\left\{h\left(\frac{j}{2J^{d}}\right),j=0,1,\cdots,2J^{d}-1\right\}\leq 2w_{f}(\sqrt{d}),\ \forall\alpha\in\mathbb{R}.$ (5.21)
Define ${\psi}_{2}(\alpha):=\tilde{\psi}_{2}(2J^{d}\alpha)$. Then, $\psi_{2}\in\mathcal{F}_{1,K}\left(16NK+30K,6\left\lceil\sqrt{2}M\right\rceil+10\right)$ and it belongs to a class of the ReLU neural network with the same width and depth as $\tilde{\psi}_{2}$. Then we have
$\left|\psi_{2}(\psi_{1}(p))-h(\psi_{1}(p))\right|=\left|\psi_{2}\left(\frac{j}{2J^{d}}\right)-h\left(\frac{j}{2J^{d}}\right)\right|=\left|\tilde{\psi}_{2}(j)-h\left(\frac{j}{2J^{d}}\right)\right|\leq w_{f}\left(\frac{\sqrt{d}}{J}\right),\ j=0,1,\cdots,2J^{d}-1.$ (5.22)
(6) We construct a function $\phi_{2}$ such that $\phi_{2}(p)$ approximates $\bar{f}(x_{p})$. Let $\phi_{2}:=\psi_{2}\circ\psi_{1}$. Note that $\psi_{1}:\mathbb{R}^{d}\to\mathbb{R}$ is a piecewise linear mapping, it follows that $\phi_{2}\in{\mathcal{F}}_{d,K}(16NK+30K,6\left\lceil\sqrt{2}M\right\rceil+10)$ from (2) in Lemma 5.7. Since $\left\lceil\sqrt{2}M\right\rceil\leq\left\lceil\frac{3}{2}M\right\rceil\leq\frac{3}{2}M+\frac{1}{2}$, we have $\mathcal{F}_{d,K}(16NK+30K,6\left\lceil\sqrt{2}M\right\rceil+10)\subseteq\mathcal{F}_{d,K}(16NK+30K,9M+13)$ from (1) in Lemma 5.7, which implies $\phi_{2}\in{\mathcal{F}}_{d,K}(16NK+30K,9M+13)$.
According to (5.22) and the construction in Part (4), for all $p\in[J]^{d}$, we have
$\left|\phi_{2}(p)-\bar{f}(x_{p})\right|=\left|\psi_{2}(\psi_{1}(p))-h(\psi_{1}(p))\right|\leq w_{f}\left(\frac{\sqrt{d}}{J}\right).$ (5.23)
Furthermore, we have from (5.21)
$0\leq\phi_{2}(x)\leq 2w_{f}(\sqrt{d}),\ \forall x\in\mathbb{R}^{d}.$ (5.24)
In summary, we define $\phi(x)=\phi_{2}\circ\phi_{1}(x)+f(\mathbf{0})-w_{f}(\sqrt{d})$. Since $\psi\in{\mathcal{F}}_{1,1}(8\left\lfloor N^{1/d}\right\rfloor+3,2\left\lfloor M^{1/d}\right\rfloor+5)$, from the expression of $\phi_{1}$ in (5.19), we have $\phi_{1}\in{\mathcal{F}}_{d,d}(8d\left\lfloor N^{1/d}\right\rfloor+3d,2M+5)$ according to (1) in Lemma 5.7 and Lemma 5.8. Note that $\phi_{2}\in{\mathcal{F}}_{d,K}(16NK+30K,9M+13)$, combined with (2) in Lemma 5.7 such that $\phi=\phi_{2}\circ\phi_{1}+f(\mathbf{0})-w_{f}(\sqrt{d})\in{\mathcal{F}}_{d,K}(\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\},11M+18)$.
From equation (5.23) and the definitions of $\phi$ and $\bar{f}$, for any $x\in[0,1]^{d}\backslash\Omega([0,1]^{d},J,\delta)$, there exists $p\in[J]^{d}$, it follows that
$\left|f(x)-\phi(x)\right|$ $=\left|\bar{f}(x)-\phi_{2}(\phi_{1}(x))\right|=\left|\bar{f}(x)-\phi_{2}(p)\right|$
$\leq\left|\bar{f}(x)-\bar{f}(x_{p})\right|+\left|\bar{f}(x_{p})-\phi_{2}(p)\right|\leq w_{f}\left(\frac{\sqrt{d}}{J}\right)+w_{f}\left(\frac{\sqrt{d}}{J}\right)=2w_{f}\left(\frac{\sqrt{d}}{J}\right),$
where the second inequality follows from $\left\|x-x_{p}\right\|_{2}\leq\sqrt{(\frac{1}{J})^{2}+\cdots+(\frac{1}{J})^{2}}=\frac{\sqrt{d}}{J}$. Note that
$J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor\geq\frac{N^{2/d}M^{2/d}\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}}{2^{5}}\geq\frac{N^{2/d}M^{2/d}\left(\log_{3}(N+2)\right)^{1/d}}{2^{6}}.$
From (5.17), we have
$\left|f(x)-\phi(x)\right|\leq 2w_{f}\left(64\sqrt{d}(N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)$
$\leq$ $2\left\lceil 64\sqrt{d}\right\rceil w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\leq 130\sqrt{d}w_{f}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right).$
For any $x\in\Omega([0,1]^{d},J,\delta)$, we have $\left|\phi(x)-f(x)\right|\leq\left|\phi_{2}\circ\phi_{1}(x)-w_{f}(\sqrt{d})\right|+\left|f(\mathbf{0})-f(x)\right|\leq 2w_{f}(\sqrt{d})$ from (5.24). ∎
Lemma 5.11.
Let $x=(x_{1},x_{2})^{\top}$, $y=(y_{1},y_{2})^{\top}$, then we have
$\left|x_{1}y_{2}-x_{2}y_{1}\right|\leq\max\left\{\left|y_{2}\right|,\left|y_{1}\right|\right\}\left\|x-y\right\|_{1}.$
Proof.
By simple calculation, we have
$\left|x_{1}y_{2}-x_{2}y_{1}\right|$ $=\left|x_{1}y_{2}-y_{1}y_{2}+y_{1}y_{2}-x_{2}y_{1}\right|\leq\left|x_{1}y_{2}-y_{1}y_{2}\right|+\left|y_{1}y_{2}-x_{2}y_{1}\right|$
$=\left|y_{2}\right|\left|x_{1}-y_{1}\right|+\left|y_{1}\right|\left|y_{2}-x_{2}\right|\leq\max\left\{\left|y_{2}\right|,\left|y_{1}\right|\right\}\left\|x-y\right\|_{1},$
which completes the proof. ∎
Theorem 5.7.
Let $\kappa$ is a continuous map from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$ and define $F=\mathrm{softmax}\circ\kappa:[0,1]^{d}\to\mathcal{Y}=\{y=[y_{1}\ \cdots\ y_{K}]^{\top}|y_{i}\geq 0\ \forall i,\sum^{K}_{i=1}y_{i}=1\}$. For any $N,M\in\mathbb{N}^{+}$, there exists $\Phi=[\phi^{(1)}\ \cdots\ \phi^{(K)}]^{\top}\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}})$ with ${\mathcal{W}}=\max\left\{8d\left\lfloor N^{1/d}\right\rfloor+3d,16NK+30K\right\}$ and ${\mathcal{D}}=11M+18$ such that the following holds
$\left\|\Phi-F\right\|_{L_{1}([0,1]^{d})}\triangleq\sum^{K}_{i=1}\left\|\phi^{(i)}-F^{(i)}\right\|_{L^{1}([0,1]^{d})}\leq 131c\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$
where $w_{\kappa}(\cdot)$ is defined in (5.16) and $c>0$ is a constant.
Proof.
From Theorem 5.6, for given continuous map $\kappa$ from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$, there exists $\phi\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}})$
$\left|\phi(x)-\kappa(x)\right|\leq\left\{\begin{array}[]{ll}130\sqrt{d}w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right),&\forall x\in[0,1]^{d}\backslash\Omega([0,1]^{d},J,\delta),\\ 2w_{\kappa}(\sqrt{d}),&\forall x\in\Omega([0,1]^{d},J,\delta).\end{array}\right.$ (5.25)
where $\mathbf{0}=[0,0,\cdots,0]^{\top}\in{\mathbb{R}}^{d}$, $J=\left\lfloor N^{1/d}\right\rfloor^{2}\left\lfloor M^{1/d}\right\rfloor^{2}\left\lfloor\left\lfloor\log_{3}(N+2)\right\rfloor^{1/d}\right\rfloor$ and $\delta$ is an arbitrary number in $(0,\frac{1}{3J}]$.
For any $x\in[0,1]^{d}$, it is clear that $\mu\left(\Omega([0,1]^{d},J,\delta)\right)\leq Jd\delta$. Then
[…]
as long that $\delta>0$ is small enough. Combining (5.25) and (5.26), we get for any $i=1,\cdots,K$
$\left\|\phi^{(i)}-\kappa^{(i)}\right\|_{L^{1}(\nu)}\leq 131\sqrt{d}w_{\kappa^{(i)}}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right).$
For any $x\in[0,1]^{d}$, set $\phi(x)=\bar{x}\in\mathbb{R}^{K}$, $\kappa(x)=\tilde{x}\in\mathbb{R}^{K}$. Then we have
$\|\mathrm{softmax}(\bar{x})-\mathrm{softmax}(\tilde{x})\|_{1}=\sum^{K}_{i=1}\left|\frac{\exp(\bar{x}_{i})}{\sum_{j=1}^{K}\exp(\bar{x}_{j})}-\frac{\exp(\tilde{x}_{i})}{\sum_{j=1}^{K}\exp(\tilde{x}_{j})}\right|$
$=$ $\sum^{K}_{i=1}\frac{\left|\exp(\bar{x}_{i})\sum_{j\neq i}\exp(\tilde{x}_{j})-\exp(\tilde{x}_{i})\sum_{j\neq i}\exp(\bar{x}_{j})\right|}{\sum_{j\neq i}\exp(\bar{x}_{j})\sum_{j\neq i}\exp(\tilde{x}_{j})}\leq e\left\|\exp(\bar{x})-\exp(\tilde{x})\right\|_{1}\leq c\left\|\bar{x}-\tilde{x}\right\|_{1},$
where the last second inequality is from Lemma 5.11, the last one is from Lagrangian mean value theorem and $c>0$ is a positive constant. Set $\Phi=\mathrm{softmax}\circ\phi$. Then
[…]
which completes the proof. ∎
Proof of Theorem 4.1..
According to Theorem 5.1 and Remark 5.2, when $\eta<\frac{K-1}{K}$, we have $\hat{f}^{\eta}_{n}=\hat{f}_{n}$ when the loss function is one of the $\ell_{1}$, CE and RCE losses. When the loss function is $\ell_{1}$, we have
$|\mathcal{L}(\hat{f}^{\eta}_{n})-\mathcal{L}(f_{0})|=$ $|\mathcal{L}(\hat{f_{n}})-\mathcal{L}(f_{0})|\leq 2\mathbb{E}_{S}\left[\sup_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left|\mathcal{L}(f)-\mathcal{L}_{n}(f)\right|\right]+\lambda\inf_{f\in\mathcal{F}_{d,K}(\mathcal{W},\mathcal{D})}\left\|f-f_{0}\right\|_{L^{1}(\nu)}$
$\leq$ $\frac{8\lambda K{\mathcal{M}}\sqrt{{\mathcal{D}}+2+\log d}}{a_{n}\sqrt{na_{n}}}+\frac{4\lambda Kn\beta_{a_{n}}}{a_{n}}+131c\sqrt{d}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/d}\right)\right\|_{1},$
where the first inequality is from Lemma 5.2, the second inequality is from Theorems 5.3 and 5.7 with
$\lambda=\left\{\begin{array}[]{ll}1,&\ell_{1}\text{\ loss},\\ \frac{1}{\epsilon},&\text{CE loss},\\ -A,&\text{RCE loss}.\end{array}\right.$
Similarly, we get Equation (4.1) from Theorem 5.5 and 5.7. ∎
6 The Curse of Dimensionality
In Subsection 5.4, the upper bound for the approximation error is given by $131c\sqrt{d}\|w_{\kappa}((N^{2}M^{2}\log_{3}(N+2))^{-1/d})\|_{1}$. However, as $d$ approaches infinity, this bound exhibits explosive growth, resulting in the curse of dimensionality. In this section, we mitigate this issue based on the following assumption.
Assumption 6.1.
The feature space is supported on a compact $s$-dimensional Riemannian manifold $\mathcal{A}$ embedded in ${\mathbb{R}}^{d}$, which can be decomposed into a union of $C$ open sets in ${\mathbb{R}}^{d}$. That is, $\mathcal{A}=\bigcup^{C}_{j=1}U_{j}$ and $C\ll d$. For each $j\in\{1,2,\dots,C\}$, there exists a bijective function $\zeta_{j}:U_{j}\to[0,1]^{s}$ with $s<d$ such that $\zeta_{j}(U_{j})=\tilde{U_{j}}$, where $\tilde{U_{j}}\subseteq[0,1]^{s}$.
Assumption 6.1 is well-founded. There is a vast amount of data generated and collected in our daily lives. Despite their diverse sources and natures, these seemingly unstructured data often possess inherent structures. One commonly used structure is that high-dimensional data may reside in a low-dimensional subspace rather than being arbitrarily distributed throughout the entire space. For instance, handwritings of the same digit with varying rotations, translations, or thickness are approximately within a low-dimensional subspace. Other examples include human face image data, where the face images of each individual under different illuminations form a low-dimensional subspace, and a video sequence with multiple moving objects, where each moving object across different frames belongs to a low-dimensional subspace.
In addition, the assumption about bijective functions is also reasonable. In the realm of compressive sensing, accurate recovery of the true sparse signal is possible when the sensing matrix fulfills the Restricted Isometry Property (RIP) [CT05, Can08], or Block Restricted Isometry Property (BRIP) [KSHT13, HLL25] condition, thereby enabling a one-to-one mapping between vectors in ${\mathbb{R}}^{d}$ and ${\mathbb{R}}^{s}$. Furthermore, the theory of differential manifold [Bra02, JSLH23] guarantees the existence of bijectives.
In the following, we will give a new estimation of the error bound of the approximation error based on Assumption 6.1.
Theorem 6.1.
Assuming that Assumption 6.1 holds, and given bijective functions $\zeta_{j}$ satisfying $\|\zeta_{j}(x_{1})-\zeta_{j}(x_{2})\|\geq Q_{2}\|x_{1}-x_{2}\|,\forall x_{1},x_{2}\in U_{j},\forall j\in\{1\cdots,C\}$, for any $N,M\in\mathbb{N}^{+}$ and any continuous map $\kappa:[0,1]^{d}\to{\mathbb{R}}^{K}$, define $F=\mathrm{softmax}\circ\kappa:[0,1]^{d}\to\mathcal{Y}$, then there exists an $f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}})$ with ${\mathcal{W}}=C\cdot\max\{8s\left\lfloor N^{1/s}\right\rfloor+3s,16NK+30K\}$ and ${\mathcal{D}}=11M+18$ such that
$\left\|f-F\right\|_{L^{1}{([0,1]^{d})}}\leq 131c\sqrt{s}Q_{1}\left\|w_{\kappa}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/s}/Q_{2}\right)\right\|_{1},$ (6.1)
where $Q_{1}>0$ are the maximal Jacobi determinants of $\zeta_{j},j=1,\cdots,C$ and $c$ is a positive constant.
Proof.
When Assumption 6.1 holds, there exist continuous bijective functions $\zeta_{j}:U_{j}\to[0,1]^{s}$ such that $\zeta_{j}(x)=\tilde{x},\forall x\in U_{j}$ and $F(x)=F\circ\zeta^{-1}_{j}(\tilde{x}),\forall x\in U_{j}$.
Define $\zeta(x)=\sum^{C}_{j=1}\zeta_{j}(x)\mathds{1}(x\in U_{j})$, then $\zeta(\cdot)$ is also a bijective function and $\zeta(x)\in[0,1]^{s},\forall x\in\cal{A}$. Set $\tilde{x}=\zeta(x)$, define $\tilde{F}(\tilde{x})=F\circ\zeta^{-1}(\tilde{x})$, then
$F(x)=F\circ\zeta^{-1}(\zeta(x))=\tilde{F}(\tilde{x})=\sum^{C}_{j=1}F\circ\zeta^{-1}_{j}(\tilde{x})\mathds{1}(\tilde{x}\in\zeta_{j}(U_{j}))$
and $\tilde{F}$ is a mappings from $[0,1]^{s}$ to $\mathcal{Y}$.
For all $j\in\{1,\cdots,C\}$, define $\tilde{F}_{j}(\tilde{x})=F\circ\zeta^{-1}_{j}(\tilde{x})\mathds{1}(\tilde{x}\in\zeta_{j}(U_{j}))=\mathrm{softmax}\circ\kappa\circ\zeta^{-1}_{j}(\tilde{x})\mathds{1}(\tilde{x}\in\zeta_{j}(U_{j}))\triangleq\mathrm{softmax}\circ\tilde{\kappa}_{j}(\tilde{x})$, then $\tilde{\kappa}:[0,1]^{s}\to{\mathbb{R}}^{K}$ is a continuous mapping. From Theorem 5.7, there exists a function $\tilde{f}_{j}\in{\mathcal{F}}_{s,K}({\mathcal{W}},{\mathcal{D}})$ for each $\tilde{F}_{j}$, with ${\mathcal{W}}=\max\{8s\left\lfloor N^{1/s}\right\rfloor+3s,16NK+30K\}$ and ${\mathcal{D}}=11M+18$ such that
$\left\|\tilde{f_{j}}-\tilde{F}_{j}\right\|_{L_{1}([0,1]^{s})}\leq 131c\sqrt{s}\sum^{K}_{i=1}w_{\tilde{\kappa}^{(i)}_{j}}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/s}\right),j=1,\cdots,C.$
Next we define
$f(x)=\sum_{j=1}^{C}\tilde{f}_{j}(\zeta(x))=\sum_{j=1}^{C}\tilde{f}_{j}\circ\zeta(x),$
then we know that $f$ is also a ReLU neural network with the same depth as $\tilde{f}_{j}$, but its width is $C$ times that of $\tilde{f}_{j}$. That is $f\in{\mathcal{F}}_{d,K}({\mathcal{W}},{\mathcal{D}})$ with ${\mathcal{W}}=C\cdot\max\left\{8s\left\lfloor N^{1/s}\right\rfloor+3s,16NK+30K\right\}$ and ${\mathcal{D}}=11M+18$. Hence
$\left\|f-F\right\|_{L^{1}{([0,1]^{d})}}=\left\|\sum_{j=1}^{C}\tilde{f}_{j}(\zeta(x))-\tilde{F}_{j}(\zeta(x))\right\|_{L^{1}{([0,1]^{d})}}\overset{\tilde{x}=\zeta(x)}{\leq}Q_{1}\left\|\sum_{j=1}^{C}\tilde{f}_{j}(\tilde{x})-\tilde{F}_{j}(\tilde{x})\right\|_{L^{1}{([0,1]^{s})}}$
$\leq$ $131\sqrt{s}Q_{1}c\sum_{i=1}^{K}w_{\tilde{\kappa}^{(i)}}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/s}\right)$
$\leq$ $131\sqrt{s}Q_{1}c\sum_{i=1}^{K}w_{\kappa^{(i)}}\left((N^{2}M^{2}\log_{3}(N+2))^{-1/s}/Q_{2}\right),$
where $Q_{1}>0$ are the maximal Jacobi determinants of $\zeta_{j},j=1,\cdots,C$. In the first inequality, $Q_{1}$ is from variable substitution $\tilde{x}=\zeta(x)$ and the last inequality is from (5.17). ∎
To prove Theorem 6.1, we construct a specific mapping $\tilde{F}$ from $[0,1]^{s}$ to ${\mathbb{R}}^{K}$, where it is defined as a sum of $C$ functions. Our proof mainly includes the following steps:
We first estimate the upper bound of the approximation error for each function. According to Theorem 5.7, for each function, we can derive an upper bound for its approximation error. This bound depends on $s$, the intrinsic dimensionality of the domain, rather than $d$.
We establish the one-to-one correspondence by bijective functions. Assumption 6.1 states that $\zeta_{j}$ for all $j$ are bijective functions. These functions establish a one-to-one correspondence between the Riemannian manifold and a transformed space. This correspondence is crucial because it allows us to map the simpler, lower-dimensional space back to the Riemannian manifold.
We estimate the upper bound for the approximation error of $F$. Given the one-to-one correspondence between $F$ and $\tilde{F}$, we can derive the upper bound for the approximation error of $F$ by analyzing the approximation error of $\tilde{F}$. Since $\tilde{F}$ lives in a lower-dimensional space (controlled by $s$), its approximation error is easier to bound and is influenced by $s$ rather than $d$.
By combining the results from the individual bounds on each function from $[0,1]^{s}$ to ${\mathbb{R}}^{K}$ and leveraging the bijective functions, we demonstrate that the overall approximation error for $F$ is controlled by a function of $s$. This is significant because it shows that, despite potentially high ambient dimensionality $d$, the complexity of approximation is determined by the lower intrinsic dimensionality $s$.
7 Conclusion
Labeling large-scale datasets is a costly and error-prone process, and even high-quality datasets are inevitably subject to mislabeling. Training deep neural networks on cheaply obtained noisy labeled datasets is a common task. In this paper, we focus on the error bounds of excess risk for classification problems with noisy labels within deep learning frameworks.
We begin with exploring loss functions with noise-tolerant properties, ensuring that the empirical minimizer on noisy data aligns with that on the true data. Next, we estimate the error bound of the excess risk, expressed as a sum of statistical error and approximation error. We assess the statistical error on a dependent (mixing) sequence, bounding it with the help of the corresponding independent block sequence. For the approximation error, we first express the classifiers as the composition of the softmax function and a continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. The main task is then to estimate the error bound for the continuous function from $[0,1]^{d}$ to ${\mathbb{R}}^{K}$. Finally, we focus on the curse of dimensionality based on the low-dimensional manifold assumption.
This work has its own limitation: it focuses solely on the $\ell_{1}$ loss, CE, and RCE loss; additional loss functions may be explored in future research.